OpenStudy (anonymous):

Solve 2x2 = 50. A. {±5} B. {±25} C. {±6} D. {±12.5}

OpenStudy (kendricklamar2014):

Step 1: Divide both sides by 2 \[\frac{ 2x^2 }{ 2 } = \frac{ 50 }{ 2 }\] Which equals: \[x^2 = 25\]

OpenStudy (kendricklamar2014):

Step 2. Take the Square Root of 25: \[x = \pm \sqrt{25}\] Which equals: \[x = 5 ~or~ x = -5\]

OpenStudy (kendricklamar2014):

\[Answer: \pm 5\]